OpenStudy (logan13):

What is 3 root of 343

OpenStudy (sweetnikkie072):

7

Directrix (directrix):

@sweetnikkie072 How did you get 7? Please show some work. Thanks.

OpenStudy (logan13):

7.5

jhonyy9 (jhonyy9):

7*7=49*7=343

OpenStudy (sweetnikkie072):

okhay..so basically you are finding cubic root of 343..honestly i just remember some basic square roots and cubic roots..what you can do is use long division too..is that what you want me to explain?

OpenStudy (sweetnikkie072):

its basically how you multiply numbers 3 times which gives you the answer..if you look at it

OpenStudy (agent0smith):

Remember learning prime factorization?  So 343 = 7*7*7 = 7^3

OpenStudy (logan13):

so 73

OpenStudy (agent0smith):

7*7*7 is not 73. Neither is 7^3. 7^3 is\[\Large 7^3\]

OpenStudy (agent0smith):

\[\Large \sqrt[3]{343} = \sqrt[3]{7^3} =?\]

OpenStudy (logan13):

so its

OpenStudy (agent0smith):

So it's...?

OpenStudy (logan13):

idk

OpenStudy (agent0smith):

Here's a hint\[\Large \sqrt[n]{x^n} = x\]

OpenStudy (logan13):

so would it be 3

OpenStudy (logan13):

@agent0smith -3

OpenStudy (agent0smith):

Look at the hint more carefully

OpenStudy (logan13):

55.5607775324

OpenStudy (logan13):

so 55

OpenStudy (sweetnikkie072):

7*7*7..so you are multiplying 7 ...three times....and you are getting 343...now my question to you is.....what is the number that you multiply 3 times to get 343?.... 7? so what is the cubic root ( 3√) of 343? 7? therefore... 3√(343)=7

OpenStudy (logan13):

so it is 7